# Interpretable Vision-Language Survival Analysis with Ordinal Inductive Bias for Computational Pathology

### v3

@@ Abstract
Histopathology Whole-Slide Images (WSIs) provide an important tool to assess cancer prognosis in computational pathology (CPATH). While existing survival analysis (SA) approaches have made exciting progress, they are generally limited to adopting highly-expressive architectures and only coarse-grained patient-level labels to learn prognostic visual representations from gigapixel WSIs. Such learning paradigm suffers from important performance bottlenecks, when facing present scarce training data and standard multi-instance learning (MIL) framework in CPATH. To overcome it, this paper, for the first time, proposes a new Vision-Language-based SA (VLSA) paradigm. Concretely, (1) VLSA is driven by pathology VL foundation models. It no longer relies on high-capability networks and shows the advantage of data efficiency. (2) In vision-end, VLSA encodes prognostic language prior and then employs it as auxiliary signals to guide the aggregating of prognostic visual features at instance level, thereby compensating for the weak supervision in MIL. Moreover, given the characteristics of SA, we propose i) ordinal survival prompt learning to transform continuous survival labels into textual prompts; and ii) ordinal incidence function as prediction target to make SA compatible with VL-based prediction. Notably, VLSA’s predictions can be interpreted intuitively by our Shapley values-based method. The extensive experiments on five datasets confirm the effectiveness of our scheme. Our VLSA could pave a new way for SA in CPATH by offering weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs. Our source code is available at https://github.com/liupei101/VLSA.
@@ 1 Introduction
However, the survival analysis (SA) of WSI data has always been faced two critical challenges. While exiting approaches have made exciting progress in overcoming these challenges, they still suffer from important performance bottlenecks due to the inherent limitations of current SA paradigm. (1) Scarce training data. Owing to many real-world factors, e.g., the difficulties in long-term patient follow-up and the concerns about patient privacy, the scale of WSI data for SA has always been limited (Lu et al., 2022; Liu et al., 2024a; Song et al., 2024), typically on the order of 1,000. Most existing SA models often overlook this and generally seek for network-level solutions to improve the performance, such as adopting highly-expressive modern networks like GNNs (Chen et al., 2021; Liu et al., 2023b; Shao et al., 2024; Wang et al., 2024) or Transformers (Huang et al., 2021; Hou et al., 2022; Liu et al., 2023a). However, when facing present small WSI data, this way is more likely to cause overfitting in prognostic models (Srivastava et al., 2014), leading to suboptimal prediction performance. (2) Learning from gigapixel images under weak supervision. Digital WSIs have extremely-high resolution, e.g., 40,000$\times$ 40,000 pixels, so each one is usually processed into a bag of multi-instances for training. With only WSI-level labels, bag-level representations are derived via a defacto weakly-supervised multi-instance learning (MIL) framework (Ilse et al., 2018; Liu et al., 2024b). It first i) learns task-specific embeddings from single instances and ii) then aggregates numerous instances (typically 10,000) into one single vector. Current SA schemes follow this, yet the whole learning process is driven entirely by patient-level labels. We argue that such paradigm could lead to inefficient representations, since SA models are only provided with overall patient-level labels while they are not only required to i) learn prognostic embeddings at a fine-grained instance level; but also to ii) select key instances from numerous candidates (Li et al., 2023).
@@
To overcome the limitations of current practices, this paper studies a new SA paradigm for CPATH, called Vision-Language Survival Analysis (VLSA). Concretely, we find that recent foundational VL models (VLMs) in CPATH, e.g., CONCH (Lu et al., 2024), offer potential means to mitigate the challenges above. First, these VLMs are pretrained on large-scale image-text pairs by task-agnostic objectives. They show surprisingly-good performance in terms of data efficiency, especially in zero-shot transfer, as highlighted in Lu et al. (2024); Javed et al. (2024). This is promising for alleviating the challenge of scarce training data. Second, VL contrastive pretraining aligns image and language in latent embedding space (Radford et al., 2021), which enables language to act as “prompt” for vision tasks. This implies that, with prior knowledge, language is likely to provide auxiliary signals to boost learning efficiency. Such additional signals could be particularly important to the weak supervision in MIL. Despite all these appealing merits, VLM-driven SA has not yet been studied. We believe there are two main reasons: i) the powerful VLMs for CPATH are developed just recently; ii) different from conventional classification, there remains a gap for SA to adapt to VLMs.
@@
Based on these insights, this paper first proposes a VLSA framework for CPATH. Different from existing VL-based schemes, VLSA comprises four core designs as follows. (1) Vision-end: it leverages language-encoded prognostic priors to guide multi-instance aggregation, producing multi-level visual presentations. (2) Language-end: considering the intrinsic ordinality in survival risks, we propose ordinal survival prompt learning to encode continuous time-to-event labels into textual prompts. (3) Prediction and optimization: To make SA compatible with VL-based prediction, we take incidence function as prediction target and introduce an ordinal inductive bias into it for regularization in optimization. (4) Prediction interpretation: with the classic Shapley values from game theory, individual prognostic risk could be interpreted from an intuitive perspective—descriptive language. The extensive experiments on five datasets verify the effectiveness of our scheme. Notably, our comparative experiments and analyses suggest that VLSA could pave a new way for SA in CPATH by offering weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs,. We summarize the main contributions of this paper as follows:
@@
1. An interpretable Vision-Language Survival Analysis framework (VLSA) is proposed for computational pathology. To our knowledge, it is the first work that studies VL-based SA in CPATH.
@@
2. Language-encoded prognostic prior is proposed to boost weakly-supervised WSI representation learning. In view of the characteristics of SA, two ordinal inductive bias terms, i.e., ordinal survival prompts and ordinal incidence function, are introduced into VLSA to enhance its performance.
@@
3. To assess model performance more rigorously, this paper conducts extensive experiments and adopts multiple metrics involving discrimination and calibration evaluation. Empirical results show that VLSA could often obtain new state-of-the-art performances with less computation costs.
@@ 2 Related Work
Survival Analysis on WSIs One key challenge of WSI survival analysis is how to effectively learn global prognostic representations from multi-instances with only patient-level survival labels. To this end, various data-driven approaches are studied. They can be roughly grouped into three categories based on the structure assumed for instances: i) cluster-based (Yao et al., 2020; Shao et al., 2021a), ii) graph-based (Chen et al., 2021; Liu et al., 2023b; Shao et al., 2024; Wang et al., 2024), and iii) sequence-based (Huang et al., 2021; Hou et al., 2022; Liu et al., 2023a). These structures are employed mainly to learn non-local prognostic embeddings at cluster, node, or global token level. However, most studies in this field currently focus more on pure vision or vision-gene representation learning (Zhou & Chen, 2023; Jaume et al., 2024). The vision-language learning paradigm, which has witnessed remarkable successes in recent years (Zhang et al., 2024), remains under-studied.
@@
Figure 1: Overview of Vision-Language Survival Analysis (VLSA). (a) WSI representation learning with language-encoded prognostic priors (Section 3.1). (b) Prediction of ordinal incidence function (Section 3.3). (c) Ordinal survival prompt learning (Section 3.2). The survival prediction of VLSA can be interpreted by quantifying each prognostic prior’s contribution to overall risk (Section 3.5).
@@ 3 Method
This section presents our VLSA, a vision-language-based survival analysis framework (Figure 1) for CPATH. We first introduce its three key parts: i) WSI representation learning with language-encoded prognostic priors (Figure 1(a), Section 3.1); ii) Ordinal survival prompt learning (Figure 1(c), Section 3.2); and iii) ordinal incidence function prediction (Figure 1(b), Section 3.3). Then, we give our overall training objectives in Section 3.4 and explain how to interpret the survival prediction of VLSA in Section 3.5. For new concepts or terminologies, this section will provide necessary explanations. Detailed elucidations could be found in Appendix B.
@@ 3.1 WSI Representation Learning with Language-encoded Prognostic Priors
Given the WSIs of one patient, we denote their processed instances by a bag ${\bm{X}}=[{\bm{x}}_{1},\dots,{\bm{x}}_{K}]^{\mathsf{T}}\in\mathbb{R}^{K\times D}$, where ${\bm{x}}_{k}\in\mathbb{R}^{D}$ represents the $k$-th instance features. Instance features are extracted by the image encoder of a foundational VLM, written as $E_{\text{image}}$. To derive WSI representations, we introduce language-encoded prognostic priors to encourage MIL models to distill valuable prognostic clues from numerous instances under weak supervision, as shown in Figure 1(a).
@@
WSI Representation In our MIL aggregator $\mathcal{A}_{\text{MIL}}$, any vector ${\bm{p}}_{m}$ is utilized as a query to aggregate key instances by cross-attention, producing one-level WSI representation (written as ${\bm{f}}_{m}\in\mathbb{R}^{D}$) corresponding to the $m$-th prognostic description ($\mathcal{T}_{\text{prog}}^{m}$). Finally, multi-level visual representations, ${\bm{F}}=[{\bm{f}}_{1},\dots,{\bm{f}}_{M}]^{\mathsf{T}}\in\mathbb{R}^{M\times D}$, are averaged and passed through a linear neural network layer to obtain the final image features ${\bm{f}}_{\text{image}}\in\mathbb{R}^{D}$. Formally, we express these steps as follows:
@@
where $\alpha$ is a fixed temperature hyper-parameter, $\text{cos}(\cdot,\cdot)$ is an operator for calculating cosine similarity, and $\text{Linear}(\cdot)$ indicates a simple linear layer with learnable weights and biases.
Justification (1) Prognostic prior knowledge: from Eq. (2), we can find that the instances whose visual features are well-aligned with the prognostic description would be selected and aggregated into one visual representation. This implies that, with the well-aligned VL embedding space provided by foundational VLMs, prognostic prior could offer additional helpful signals and compensate for the weak supervision in MIL. (2) Multi-level visual representations: they are designed to provide sufficient prognostic clues because there are usually more than one prognostic markers in WSIs, e.g., nuclear atypia, perineural invasion, mitotic activity, etc. Moreover, multi-level prognostic clues could enable us to decouple and interpret individual survival predictions from intuitive perspectives (i.e., descriptive language). Our interpretation method is elaborated in Section 3.5.
@@ 3.2 Ordinal Survival Prompt Learning
Different from classification, SA provides continuous time-to-event labels, i.e., ${\mathbb{Y}}=\{t,\delta\}$, where $t$ is the last follow-up time and $\delta\in\{0,1\}$ is an event indicator at $t$. Due to this intrinsic difference, obtaining SA prompts would be more challenging compared with VL-based classification. To address it, we propose ordinal survival prompt learning that encodes time-to-event labels into textual survival prompts for VL-based prediction, as depicted in Figure 1(c).
Time Discretization Following the convention of discrete-time SA (Haider et al., 2020), we first uniformly discretize time into a set of non-overlapping bins, $[T_{0},T_{1}),[T_{1},T_{2}),\cdots,[T_{C-1},T_{C})$. For any $t\in[T_{c-1},T_{c})$, we assign its time-discrete label $c$, where $c\in[1,C]$. These time bins are equal in length; $C$ is determined by $\sqrt{N_{e}}$, where $N_{e}$ is the number of patients with event in survival data. This setting often leads to better calibration for discrete SA models (see Appendix E).
@@
where ${\bm{D}}_{c,b}$ is an element of matrix ${\bm{D}}\in\mathbb{R}^{C\times B}$ representing the ordering distance between the $c$-th class prompt and the $b$-th base class prompt, and $W(\cdot)$ is a function determining the interpolation weight based on distance. A closer distance usually indicates a greater interpolation weight. ${\bm{D}}_{c,b}$ could be simply set by ${\bm{D}}_{c,b}=|(c-1)-(b-1)\cdot(C-1)/(B-1)|$. For linear interpolation, $W({\bm{D}}_{c,b})=1-\frac{{\bm{D}}_{c,b}}{C-1}$. (3) For survival prompts, each class prompt is concatenated with ${\bm{V}}_{\text{ctx}}$ at its token level, followed by passing through frozen $E_{\text{text}}$ to produce the final ordinal survival prompts, denoted as ${\bm{F}}_{\text{text}}=[{\bm{f}}_{\text{text}}^{1},\dots,{\bm{f}}_{\text{text}}^{C}]^{\mathsf{T}}\in\mathbb{R}^{C\times D}$. Namely, there is
@@
Justification (1) Ordinal inductive bias: in common sense, as the length of survival time decreases, the corresponding risk of death gradually becomes high. This implies the ordinality of survival classes; such inductive bias could be considered in designing prompts for SA. More explanations and discussions could be found in Appendix C. (2) Class prompt interpolation: on one hand, most textual encoders are lacking in the sensitivity to numbers’ ordinality (Thawani et al., 2021; Paiss et al., 2023), so it may be unsuitable to directly adopt numerical classes as textual prompts. On the other hand, when the number of classes becomes large, e.g., $C=10$, it would be intractable to manually design fine-grained prognosis risk descriptions at $C$ different levels. In view of these obstacles, we adopt a few base prompts ($\lambda\leq 4$) and then utilize an interpolation-based strategy (Li et al., 2022) to preserve ordinalitity in survival prompts.

### v4

@@ Abstract
Histopathology Whole-Slide Images (WSIs) provide an important tool to assess cancer prognosis in computational pathology (CPATH). While existing survival analysis (SA) approaches have made exciting progress, they are generally limited to adopting highly-expressive network architectures and only coarse-grained patient-level labels to learn visual prognostic representations from gigapixel WSIs. Such learning paradigm suffers from critical performance bottlenecks, when facing present scarce training data and standard multi-instance learning (MIL) framework in CPATH. To overcome it, this paper, for the first time, proposes a new Vision-Language-based SA (VLSA) paradigm. Concretely, (1) VLSA is driven by pathology VL foundation models. It no longer relies on high-capability networks and shows the advantage of data efficiency. (2) In vision-end, VLSA encodes textual prognostic prior and then employs it as auxiliary signals to guide the aggregating of visual prognostic features at instance level, thereby compensating for the weak supervision in MIL. Moreover, given the characteristics of SA, we propose i) ordinal survival prompt learning to transform continuous survival labels into textual prompts; and ii) ordinal incidence function as prediction target to make SA compatible with VL-based prediction. Notably, VLSA’s predictions can be interpreted intuitively by our Shapley values-based method. The extensive experiments on five datasets confirm the effectiveness of our scheme. Our VLSA could pave a new way for SA in CPATH by offering weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs. Our source code is available at https://github.com/liupei101/VLSA.
@@ 1 Introduction
However, the survival analysis (SA) of WSI data has always faced two critical challenges. While existing approaches have made exciting progress in overcoming these challenges, they still suffer from performance bottlenecks due to the inherent limitations of current SA paradigm. (1) Scarce training data. Owing to many real-world factors, e.g., the difficulties in long-term patient follow-up and the concerns about patient privacy, the scale of WSI data for SA has always been limited (Lu et al., 2022; Liu et al., 2024a; Song et al., 2024a), typically on the order of 1,000. Most existing SA models often overlook this and generally seek for network-level solutions to improve the performance, such as adopting highly-expressive modern networks like GNNs (Chen et al., 2021a; Liu et al., 2023b; Shao et al., 2024; Wang et al., 2024) or Transformers (Huang et al., 2021; Hou et al., 2022; Liu et al., 2023a). However, when facing present small WSI data, this way is more likely to cause overfitting in deep learning models (Srivastava et al., 2014), leading to suboptimal prediction performance. (2) Learning from gigapixel images under weak supervision. Digital WSIs have extremely-high resolution, e.g., 40,000$\times$ 40,000 pixels, so each one is usually processed into a bag of multi-instances for training. With only WSI-level labels, bag-level representations are derived via a defacto weakly-supervised multi-instance learning (MIL) framework (Ilse et al., 2018; Liu et al., 2024b). It first i) learns task-specific embeddings from single instances and ii) then aggregates numerous instances (typically 10,000) into one single vector. Current SA schemes follow this way, yet the whole learning process is driven entirely by patient-level labels. We argue that such paradigm could lead to inefficient representations, since SA models are only provided with overall patient-level labels while they are not only required to i) learn prognostic embeddings at a fine-grained instance level; but also to ii) select key instances from numerous candidates (Li et al., 2023).
@@
To overcome the limitations of current practices, this paper studies a new SA paradigm for CPATH, called Vision-Language Survival Analysis (VLSA). Concretely, we find that recent VL foundation models (VLMs) in CPATH, e.g., CONCH (Lu et al., 2024), offer potential means to mitigate the challenges above. First, these VLMs are pretrained on large-scale image-text pairs by task-agnostic objectives. They show surprisingly-good performance in terms of data efficiency, especially in zero-shot transfer, as highlighted in Lu et al. (2024); Javed et al. (2024). This is promising for alleviating the challenge of scarce training data. Second, VL contrastive pretraining aligns image and language in latent embedding space (Radford et al., 2021), which enables language to act as “prompt” for vision tasks. This implies that, with prior knowledge, language is likely to provide auxiliary signals to boost learning efficiency. Such additional signals could be particularly helpful to improve the weak supervision in MIL. Despite all these appealing merits, VLM-driven SA has not yet been studied. We believe there are two main reasons: i) the powerful VLMs for CPATH are developed just recently; ii) different from classification, there remains a gap for SA to be adapted to VLMs.
@@
Based on these insights, this paper first proposes a VLSA framework for CPATH. Different from existing VL-based schemes, VLSA comprises four core designs as follows. (1) Vision-end: it leverages language-encoded prognostic priors to guide multi-instance aggregation, producing multi-level visual presentations. (2) Language-end: considering the intrinsic ordinality in survival risks, we propose ordinal survival prompt learning to encode continuous time-to-event labels into textual prompts. (3) Prediction and optimization: To make SA compatible with VL-based prediction, we take incidence function as prediction target and introduce an ordinal inductive bias into it for regularization in optimization. (4) Prediction interpretation: with the classic Shapley values from game theory, individual prognostic risk could be interpreted from an intuitive perspective—descriptive language. The extensive experiments on five datasets verify the effectiveness of our scheme. Notably, our comparative experiments and analyses suggest that VLSA could pave a new way for SA in CPATH by offering weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs. We summarize the main contributions of this paper as follows:
@@
$\bullet$ An interpretable Vision-Language Survival Analysis framework (VLSA) is proposed for computational pathology. To our knowledge, it is the first work that studies VL-based SA in CPATH.
@@
$\bullet$ Language-encoded prognostic prior is proposed to boost weakly-supervised WSI representation learning. In view of the characteristics of SA, two ordinal inductive bias terms, i.e., ordinal survival prompts and ordinal incidence function, are introduced into VLSA to enhance the performance.
@@
$\bullet$ To assess model performance more rigorously, this paper conducts extensive experiments and adopts multiple metrics for discrimination and calibration evaluation. Empirical results show that VLSA could often obtain new state-of-the-art performances with less computation costs.
@@ 2 Related Work
Survival Analysis on WSIs One key challenge of WSI survival analysis is effectively learning global prognostic representations from multi-instances with only patient-level survival labels. To this end, various data-driven approaches are studied. They can be roughly grouped into three categories based on the structure assumed for instances: i) cluster-based (Yao et al., 2020; Shao et al., 2021a), ii) graph-based (Chen et al., 2021a; Liu et al., 2023b; Shao et al., 2024; Wang et al., 2024), and iii) sequence-based (Huang et al., 2021; Hou et al., 2022; Liu et al., 2023a). These structures are employed to learn non-local prognostic embeddings at cluster, node, or global token level. However, most studies in this field currently focus more on pure vision or vision-gene representation learning (Zhou & Chen, 2023; Jaume et al., 2024). The vision-language learning paradigm, which has witnessed remarkable successes in recent years (Zhang et al., 2024a), remains under-studied.
@@
Figure 1: Overview of Vision-Language Survival Analysis (VLSA). (a) WSI representation learning with language-encoded prognostic priors (Section 3.1). (b) Ordinal survival prompt learning (Section 3.2). (c) Prediction of ordinal incidence function (Section 3.3). The survival prediction of VLSA can be interpreted by quantifying each prognostic prior’s contribution to risk (Section 3.5).
@@ 3 Method
This section presents our VLSA, a vision-language-based survival analysis framework (Figure 1) for CPATH. We first introduce its three key parts: i) WSI representation learning with language-encoded prognostic priors (Figure 1(a), Section 3.1); ii) Ordinal survival prompt learning (Figure 1(b), Section 3.2); and iii) ordinal incidence function prediction (Figure 1(c), Section 3.3). Then, we give our overall training objectives in Section 3.4 and explain how to interpret the survival prediction of VLSA in Section 3.5. For new concepts or terminologies, this section will provide concise explanations. Detailed elucidations could be found in Appendix A.
@@ 3.1 WSI Representation Learning with Language-encoded Prognostic Priors
Given the WSIs of one patient, we denote their processed instances by a bag ${\bm{X}}=[{\bm{x}}_{1},\dots,{\bm{x}}_{K}]^{\mathsf{T}}\in\mathbb{R}^{K\times D}$, where ${\bm{x}}_{k}\in\mathbb{R}^{D}$ represents the $k$-th instance feature vector. Instance features are extracted by the image encoder of a VLM, written as $E_{\text{image}}$. To derive WSI representations, we introduce language-encoded prognostic priors to encourage MIL models to distill valuable prognostic clues from numerous instances under weak supervision, as shown in Figure 1(a).
@@
WSI Representation In the MIL aggregator $\mathcal{A}_{\text{MIL}}$, any prior ${\bm{p}}_{m}$ is utilized as a query to aggregate key instances by similarity matching and weighted averaging, producing one-level WSI representation (${\bm{f}}_{m}\in\mathbb{R}^{D}$) corresponding to the $m$-th prognostic description ($\mathcal{T}_{\text{prog}}^{m}$). Finally, multi-level visual representations, ${\bm{F}}=[{\bm{f}}_{1},\dots,{\bm{f}}_{M}]^{\mathsf{T}}\in\mathbb{R}^{M\times D}$, are averaged and passed through a linear neural network layer to obtain the final WSI representation ${\bm{f}}_{\text{image}}\in\mathbb{R}^{D}$. In summary, there are
@@
where $\alpha$ is a fixed temperature hyper-parameter, $\text{cos}(\cdot,\cdot)$ is an operator that calculates cosine similarity, and $\text{Linear}(\cdot)$ indicates a simple linear layer with learnable weights and biases.
Justification (1) Prognostic prior knowledge: from Eq. (2), we can find that, the instances whose visual features are well-aligned with the prognostic description, would be selected and aggregated into one visual representation. This implies that, with the well-aligned VL embedding space provided by VLMs, prognostic prior could offer additional helpful signals and compensate for the inherent weak supervision in MIL. (2) Multi-level visual representations: they are designed to provide sufficient prognostic clues because there are usually more than one prognostic markers in WSIs, e.g., nuclear atypia, perineural invasion, mitotic activity, etc. Moreover, multi-level prognostic clues could enable us to decouple and interpret individual survival predictions from intuitive perspectives (i.e., descriptive language). Our interpretation method is elaborated in Section 3.5. (3) Difference from multi-modal representation learning: our textual description of prognostic prior is utilized to provide the weight of each instance for MIL aggregation and mainly plays a guidance role. The final representation in vision-end comes from WSI features, not a multi-modal representation fusing text and vision features. This is different from current multi-modal SA models (Chen et al., 2021b; Xu & Chen, 2023; Zhou & Chen, 2023; Xiong et al., 2024; Song et al., 2024b).
@@ 3.2 Ordinal Survival Prompt Learning
Different from classification, SA provides continuous time-to-event labels, i.e., ${\mathbb{Y}}=\{t,\delta\}$, where $t$ is the last follow-up time and $\delta\in\{0,1\}$ is an event indicator at $t$. Due to this intrinsic difference, obtaining SA prompts would be more challenging, compared with VL-based classification. To address it, we propose ordinal survival prompt learning that encodes time-to-event labels into textual survival prompts for VL-based prediction, as depicted in Figure 1(b).
Time Discretization Following the convention of discrete-time SA (Haider et al., 2020), we first uniformly discretize time into a set of non-overlapping bins, $[T_{0},T_{1}),[T_{1},T_{2}),\cdots,[T_{C-1},T_{C})$. For any $t\in[T_{c-1},T_{c})$, we assign its corresponding time-discrete label $c$, where $c\in[1,C]$. These time bins are equal in length; $C$ is determined by $\sqrt{N_{e}}$, where $N_{e}$ is the number of patients with $\delta=1$. Refer to Appendix C.5 for more details and Appendix D.6 for experiments on these settings.
@@
where ${\bm{D}}_{c,b}$ is an element of matrix ${\bm{D}}\in\mathbb{R}^{C\times B}$ representing the ordering distance between the $c$-th class prompt and the $b$-th base class prompt, and $W(\cdot)$ is a function determining the interpolation weight based on distance. A closer distance usually indicates a greater interpolation weight. ${\bm{D}}_{c,b}$ could be simply set by ${\bm{D}}_{c,b}=|(c-1)-(b-1)\cdot(C-1)/(B-1)|$. For linear interpolation, $W({\bm{D}}_{c,b})=1-\frac{{\bm{D}}_{c,b}}{C-1}$. (3) For survival prompts, each class prompt is concatenated with ${\bm{V}}_{\text{ctx}}$ at the token level, followed by passing through frozen $E_{\text{text}}$ to produce the final ordinal survival prompts, denoted as ${\bm{F}}_{\text{text}}=[{\bm{f}}_{\text{text}}^{1},\dots,{\bm{f}}_{\text{text}}^{C}]^{\mathsf{T}}\in\mathbb{R}^{C\times D}$. Namely, there is
@@
Justification (1) Ordinal inductive bias: in common sense, as the length of survival time decreases, the corresponding death risk becomes increasingly high. This implies the ordinality between survival classes; such inductive bias could be considered when designing prompts for SA. More explanations and discussions could be found in Appendix B. (2) Class prompt interpolation: on one hand, most textual encoders are lacking in the sensitivity to numbers’ ordinality (Thawani et al., 2021; Paiss et al., 2023), so it may be unsuitable to directly adopt numerical classes as textual prompts. On the other hand, when the number of classes becomes large, e.g., $C=10$, it would be intractable to manually design fine-grained prognosis risk descriptions at $C$ different levels. In view of these obstacles, we adopt a few base prompts ($\lambda\leq 4$) and then utilize an interpolation-based strategy (Li et al., 2022) to preserve the ordinality between survival prompts.